# Fewer Hallucinations, More Verification: A Three-Stage LLM-Based Framework for ASR Error Correction

### v2

@@ Abstract
Index Terms:
automatic speech recognition, error correction, chain of thought, large language models
@@ I Introduction
Text error correction has become an essential component in modern speech recognition systems, aiming to reduce word error rates (WER) [1, 2, 3, 4]. Early methods typically relied on shallow neural networks or rule-based approaches to handle simple error patterns [5, 6]. Recent advances in deep learning have led to two dominant paradigms: autoregressive sequence-to-sequence models [7, 8, 9] and non-autoregressive edit-based models [1, 2]. Autoregressive models exploit encoder-decoder architectures with Connectionist Temporal Classification (CTC) loss [10], including translation-style correction frameworks [7, 8] and entity-aware transformers [9]. In parallel, non-autoregressive edit-based models such as FastCorrect [1, 2] and SoftCorrect [4] predict edit operations through duration modeling and integrate CTC loss with sequence-to-sequence frameworks to enhance error detection.
While conventional approaches depend heavily on large labeled datasets and task-specific training, the emergence of large language models (LLMs) offers new opportunities due to their strong contextual adaptability. LLMs have demonstrated remarkable performance across a wide range of language-related tasks, including text classification [11], information extraction [12], and dialogue systems [13]. This raises a fundamental question: Can a general-purpose LLM correct ASR errors directly without fine-tuning or additional information?
Current LLM applications typically follow two paradigms: prompt-based methods and fine-tuning strategies. Prompt-based methods aim to stimulate latent capabilities, employing techniques such as generative prompting for commonsense reasoning [14] and Chain-of-Thought (CoT) prompting to enhance reasoning [15]. Fine-tuning approaches, often using low-rank adaptation (LoRA) [16], improve task-specific performance but remain resource-intensive. Building on the aforementioned studies, recent works have explored LLM integration into ASR error correction pipelines [17, 18, 19], achieving notable CER/WER reductions. However, these methods typically require either domain-specific fine-tuning or additional contextual information, which constrains their scalability in real-world applications. Therefore, leveraging the general knowledge embedded in LLMs without fine-tuning or external inputs emerges as a more practical and scalable alternative. Nevertheless, directly applying general LLMs often results in hallucination issues [20], posing a significant challenge to reliable correction.
@@ II-A LLM-based ASR Error Correction
In recent years, integrating LLMs into ASR error correction pipelines has attracted increasing attention. Min and Wang [20] investigated the direct application of LLMs for error correction and concluded that models such as GPT-4o are ineffective due to hallucination issues. Ma et al. [19] explored combining N-best rescoring with LLM-based correction; however, obtaining the N-best list from the ASR system may impose additional costs in practice. Yang et al. [18] compared LLM-based rescoring and generation methods, demonstrating that the latter outperforms the former when domain-specific information is provided or the LLM is fine-tuned. Nevertheless, such approaches require either domain knowledge or fine-tuning.
@@ II-B A Potential Catastrophe: Hallucination Risks in LLMs
Hallucinations in LLMs refer to instances where the generated responses, while grammatically correct, fluent, and plausible, deviate from the input or contradict factual information [21, 22]. In the context of ASR error correction, model-induced hallucinations are closely linked to transcription errors. Min and Wang [20] conducted an empirical study on directly integrating LLMs into ASR systems, confirming that powerful generative models can inadvertently introduce hallucination errors by fabricating content absent from the original audio.
@@ III-A Reliable LLM Correction Framework
The overall framework of RLLM-CF is illustrated in Figure 1. It consists of three main components: Error Pre-Detection, Chain-of-Thought Subtask Iterative Correction, and Reasoning Process Verification. The prompt design, illustrated in Figure III-D, incorporates three additional few-shot correction examples demonstrating specific reasoning steps, which are omitted for clarity. The complete correction algorithm is outlined in Algorithm 1.
@@ 11 Stage 2: Iterative Correction;
16       $\mathcal{C}(y^{*},y)=\begin{cases}1,&\text{if LLM has more confidence in }y^{*}\text{ than }y\\ 0,&\text{otherwise}\end{cases}$;
@@ III-D Answer Verification
Following the correction, a verification step is conducted to ensure compliance with task instructions. Specifically, we employ the model’s self-verification feedback to assess: (1) whether the answer conforms to the required format, and (2) whether all reasoning steps are correctly completed. Only when both criteria are satisfied is the corrected output accepted; otherwise, the original input is retained.
@@ IV Experiment
The experimental section is organized as follows. Section IV-A investigates and validates LLM hallucination phenomena. Section IV-B and Section IV-C describe the experimental setup, including dataset selection and model configurations. Experimental results on AISHELL-1 [23], AISHELL-2 [24], and LibriSpeech [25] are reported in Section IV-D. Finally, Section IV-E presents an overall performance analysis, including a detailed evaluation of noun recall rates.
@@ IV-B The Dataset Configuration
Attention[26] 4.65 0.29 0.12 - 5.06 79.32
Attention Rescore[27] 4.42 0.12 0.08 - 4.62 79.93
CTC Greedy Search[10] 4.95 0.13 0.09 - 5.17 77.91
CTC Prefix Search[10] 4.95 0.13 0.09 - 5.17 77.88
Attention[26] GPT-4o 3.67 0.44 0.21 -0.74(14.6%) 4.32 82.11
Attention Rescore[27] 3.45 0.34 0.22 -0.61(13%) 4.01 82.96
CTC Greedy Search [10] 3.56 0.32 0.18 -1.11(21%) 4.06 82.24
CTC Prefix Search[10] 3.56 0.30 0.20 -1.11(21%) 4.06 81.95
Attention[26] Deepseek v2 4.00 0.49 0.20 -0.37(7.3%) 4.69 81.05
Attention Rescore[27] 3.73 0.30 0.18 -0.41(8.8%) 4.21 81.98
CTC Greedy Search[10] 3.96 0.35 0.20 -0.66(12.7%) 4.51 80.62
CTC Prefix Search[10] 3.95 0.36 0.17 -0.69(13%) 4.48 80.46
@@
Attention[26] 2.71/6.68 0.72/1.14 0.40/0.96 - 3.82/8.79 94.76/88.20
Attention Rescore[27] 2.72/7.05 0.28/0.81 0.36/0.90 - 3.35/8.77 95.00/87.91
CTC Greedy Search[10] 3.00/7.59 0.33/0.99 0.37/0.93 - 3.77/9.52 94.40/86.93
CTC Prefix Search[10] 3.05/7.58 0.32/0.98 0.38/0.95 - 3.75/9.50 94.42/87.05
Attention[26] GPT-4o 2.43/6.00 0.73/1.19 0.47/1.12 -0.18/-0.47 3.64/8.32 95.67/90.40
Attention Rescore[27] 2.47/6.39 0.25/0.73 0.46/1.26 -0.16/-0.39 3.19/8.38 96.05/90.10
CTC Greedy Search[10] 2.63/6.36 0.31/0.88 0.49/1.20 -0.34/-1.07 3.43/8.45 95.57/90.08
CTC Prefix Search[10] 2.75/6.30 0.32/0.89 0.51/1.22 -0.17/-1.09 3.58/8.41 95.55/89.82
@@
We select three datasets as evaluation benchmarks, including the Chinese datasets AISHELL-1 [23] and AISHELL-2 [24], and the English dataset LibriSpeech [25], aiming to evaluate performance on both Chinese and English data. The LibriSpeech test set is partitioned into two subsets: test-clean, containing high-quality audio, and test-other, comprising noisier audio samples.
@@ IV-C The Model Configuration
In our experiment, we adopt a conformer-based attention-encoder-decoder (AED) ASR model trained with the WeNet toolkit [28], following the U2++[29] method, across three datasets. For AISHELL-1 and LibriSpeech, a non-streaming model architecture is employed, and four decoding strategies are evaluated: Attention, Attention Rescoring, CTC Greedy Search, and CTC Prefix Search.
@@
For LLMs, we select GPT-4o [30] and DeepSeek-V2 [31], both demonstrating strong multilingual capabilities in Chinese and English. The decoding temperature is set to 0.2, and the top-$p$ parameter is set to 0.8.
@@ IV-D Experiment Results
On the LibriSpeech test-clean subset, GPT-4o achieved the lowest WER of 3.19%, corresponding to an average reduction of 0.21 percentage points (5.72% relative reduction) and a maximum reduction of 0.34 percentage points (9%). On the test-other subset, a WER of 8.32% was achieved, with a maximum relative reduction of 1.09 percentage points (11.4%).
@@ IV-E Result Analysis
To analyze noun recall within substitution errors, noun filtering was performed using the method proposed in [32] for Chinese (AISHELL-1/2) and [33] for English (LibriSpeech). The analysis reveals an improvement in noun recall following correction.
@@ V-A Framework Analysis
To evaluate the effectiveness of each framework component, a component-wise ablation study was conducted on AISHELL-1 using DeepSeek-V2. Results are summarized in Table V. Consistent with prior findings [20], using LLMs with simple prompting for error correction often leads to increased error rates, as reflected by the baseline results.
@@
By integrating iterative reasoning into the framework, the CER was reduced to 4.80%. Finally, implementing answer verification further reduced the CER to 4.67%, effectively mitigating hallucination-induced errors in LLM corrections.
@@ V-C Token Consumption Analysis
To reduce input token consumption during experiments, multiple sentences were grouped together for inference. Empirical observations suggest that this strategy only slightly degrades performance. The specific token consumption is summarized in Table V-A.

### v3

@@ I Introduction
Text error correction has become an essential component in modern speech recognition systems, aiming to reduce word error rates (WER) [lengFastCorrectFastError2021, lengFastCorrect2Fast2022, shenMaskCorrectTokens2022, lengSoftCorrectErrorCorrection2023]. Early methods typically relied on shallow neural networks or rule-based approaches to handle simple error patterns [anantaramRepairingASROutput2018, shivakumarLearningMistakesImproving2019]. Recent advances in deep learning have led to two dominant paradigms: autoregressive sequence-to-sequence models [maniASRErrorCorrection2020, dharoAutomaticCorrectionASR2016, wangASRErrorCorrection2020] and non-autoregressive edit-based models [lengFastCorrectFastError2021, lengFastCorrect2Fast2022]. Autoregressive models exploit encoder-decoder architectures with Connectionist Temporal Classification (CTC) loss [gravesConnectionistTemporalClassification2006], including translation-style correction frameworks [maniASRErrorCorrection2020, dharoAutomaticCorrectionASR2016] and entity-aware transformers [wangASRErrorCorrection2020]. In parallel, non-autoregressive edit-based models such as FastCorrect [lengFastCorrectFastError2021, lengFastCorrect2Fast2022] and SoftCorrect [lengSoftCorrectErrorCorrection2023] predict edit operations through duration modeling and integrate CTC loss with sequence-to-sequence frameworks to enhance error detection.
While conventional approaches depend heavily on large labeled datasets and task-specific training, the emergence of large language models (LLMs) offers new opportunities due to their strong contextual adaptability. LLMs have demonstrated remarkable performance across a wide range of language-related tasks, including text classification [howard2018universal], information extraction [pai2024survey], and dialogue systems [bae2022buildingAR]. This raises a fundamental question: Can a general-purpose LLM correct ASR errors directly without fine-tuning or additional information?
Current LLM applications typically follow two paradigms: prompt-based methods and fine-tuning strategies. Prompt-based methods aim to stimulate latent capabilities, employing techniques such as generative prompting for commonsense reasoning [liuGeneratedKnowledgePrompting2022] and Chain-of-Thought (CoT) prompting to enhance reasoning [weiChainthoughtPromptingElicits2023]. Fine-tuning approaches, often using low-rank adaptation (LoRA) [huLoRALowRankAdaptation2021], improve task-specific performance but remain resource-intensive. Building on the aforementioned studies, recent works have explored LLM integration into ASR error correction pipelines [ma2025asr, yang2023generative, ma2023generativelargelanguagemodels], achieving notable CER/WER reductions. However, these methods typically require either domain-specific fine-tuning or additional contextual information, which constrains their scalability in real-world applications. Therefore, leveraging the general knowledge embedded in LLMs without fine-tuning or external inputs emerges as a more practical and scalable alternative. Nevertheless, directly applying general LLMs often results in hallucination issues [minExploringIntegrationLarge2024], posing a significant challenge to reliable correction.
@@ II-A LLM-based ASR Error Correction
In recent years, integrating LLMs into ASR error correction pipelines has attracted increasing attention. Min and Wang [minExploringIntegrationLarge2024] investigated the direct application of LLMs for error correction and concluded that models such as GPT-4o are ineffective due to hallucination issues. Ma et al. [ma2023generativelargelanguagemodels] explored combining N-best rescoring with LLM-based correction; however, obtaining the N-best list from the ASR system may impose additional costs in practice. Yang et al. [yang2023generative] compared LLM-based rescoring and generation methods, demonstrating that the latter outperforms the former when domain-specific information is provided or the LLM is fine-tuned. Nevertheless, such approaches require either domain knowledge or fine-tuning.
@@ II-B A Potential Catastrophe: Hallucination Risks in LLMs
Hallucinations in LLMs refer to instances where the generated responses, while grammatically correct, fluent, and plausible, deviate from the input or contradict factual information [huang2023survey, bai2024hallucination]. In the context of ASR error correction, model-induced hallucinations are closely linked to transcription errors. Min and Wang [minExploringIntegrationLarge2024] conducted an empirical study on directly integrating LLMs into ASR systems, confirming that powerful generative models can inadvertently introduce hallucination errors by fabricating content absent from the original audio.
@@ III-A Reliable LLM Correction Framework
The overall framework of RLLM-CF is illustrated in Figure 1. It consists of three main components: Error Pre-Detection, Chain-of-Thought Subtask Iterative Correction, and Reasoning Process Verification. The prompt design, illustrated in Figure III-D, incorporates three additional few-shot correction examples demonstrating specific reasoning steps, which are omitted for clarity. For Chinese error correction, we use the translated version of this prompt. The complete correction algorithm is outlined in Algorithm 1.
@@ 11 Stage 2: Iterative Correction;
[…]
@@ III-D Answer Verification
Following the correction, a verification step is conducted to ensure compliance with task instructions. Specifically, we employ the model’s output to assess: (1) whether the answer conforms to the required format, and (2) whether all reasoning steps are correctly completed. Only when both criteria are satisfied is the corrected output accepted; otherwise, the original input is retained.
@@ IV Experiment
The experimental section is organized as follows. Section IV-A investigates and validates LLM hallucination phenomena. Section IV-B and Section IV-C describe the experimental setup, including dataset selection and model configurations. Experimental results on AISHELL-1 [buAISHELL1OpensourceMandarin2017], AISHELL-2 [duAISHELL2TransformingMandarin2018], and LibriSpeech [panayotovLibrispeechASRCorpus2015] are reported in Section IV-D. Finally, Section IV-E presents an overall performance analysis, including a detailed evaluation of noun recall rates.
@@ IV-B The Dataset Configuration
Attention[chan2016listen] 4.65 0.29 0.12 - 5.06 79.32
Attention Rescore[zhang2020unified] 4.42 0.12 0.08 - 4.62 79.93
CTC Greedy Search[gravesConnectionistTemporalClassification2006] 4.95 0.13 0.09 - 5.17 77.91
CTC Prefix Search[gravesConnectionistTemporalClassification2006] 4.95 0.13 0.09 - 5.17 77.88
Attention[chan2016listen] GPT-4o 3.67 0.44 0.21 -0.74(14.6%) 4.32 82.11
Attention Rescore[zhang2020unified] 3.45 0.34 0.22 -0.61(13%) 4.01 82.96
CTC Greedy Search [gravesConnectionistTemporalClassification2006] 3.56 0.32 0.18 -1.11(21%) 4.06 82.24
CTC Prefix Search[gravesConnectionistTemporalClassification2006] 3.56 0.30 0.20 -1.11(21%) 4.06 81.95
Attention[chan2016listen] Deepseek v2 4.00 0.49 0.20 -0.37(7.3%) 4.69 81.05
Attention Rescore[zhang2020unified] 3.73 0.30 0.18 -0.41(8.8%) 4.21 81.98
CTC Greedy Search[gravesConnectionistTemporalClassification2006] 3.96 0.35 0.20 -0.66(12.7%) 4.51 80.62
CTC Prefix Search[gravesConnectionistTemporalClassification2006] 3.95 0.36 0.17 -0.69(13%) 4.48 80.46
@@
Attention[chan2016listen] 2.71/6.68 0.72/1.14 0.40/0.96 - 3.82/8.79 94.76/88.20
Attention Rescore[zhang2020unified] 2.72/7.05 0.28/0.81 0.36/0.90 - 3.35/8.77 95.00/87.91
CTC Greedy Search[gravesConnectionistTemporalClassification2006] 3.00/7.59 0.33/0.99 0.37/0.93 - 3.77/9.52 94.40/86.93
CTC Prefix Search[gravesConnectionistTemporalClassification2006] 3.05/7.58 0.32/0.98 0.38/0.95 - 3.75/9.50 94.42/87.05
Attention[chan2016listen] GPT-4o 2.43/6.00 0.73/1.19 0.47/1.12 -0.18/-0.47 3.64/8.32 95.67/90.40
Attention Rescore[zhang2020unified] 2.47/6.39 0.25/0.73 0.46/1.26 -0.16/-0.39 3.19/8.38 96.05/90.10
CTC Greedy Search[gravesConnectionistTemporalClassification2006] 2.63/6.36 0.31/0.88 0.49/1.20 -0.34/-1.07 3.43/8.45 95.57/90.08
CTC Prefix Search[gravesConnectionistTemporalClassification2006] 2.75/6.30 0.32/0.89 0.51/1.22 -0.17/-1.09 3.58/8.41 95.55/89.82
@@
We select three datasets as evaluation benchmarks, including the Chinese datasets AISHELL-1 [buAISHELL1OpensourceMandarin2017] and AISHELL-2 [duAISHELL2TransformingMandarin2018], and the English dataset LibriSpeech [panayotovLibrispeechASRCorpus2015], aiming to evaluate performance on both Chinese and English data. The LibriSpeech test set is partitioned into two subsets: test-clean, containing high-quality audio, and test-other, comprising noisier audio samples.
@@ IV-C The Model Configuration
In our experiment, we adopt a conformer-based attention-encoder-decoder (AED) ASR model trained with the WeNet toolkit [yaoWeNetProductionOriented2021], following the U2++[wu2106u2++] method, across three datasets. For AISHELL-1 and LibriSpeech, a non-streaming model architecture is employed, and four decoding strategies are evaluated: Attention, Attention Rescoring, CTC Greedy Search, and CTC Prefix Search.
@@
For LLMs, we select GPT-4o [openaiGPT4TechnicalReport2024] and DeepSeek-V2 [deepseek-aiDeepseekv2StrongEconomical2024], both demonstrating strong multilingual capabilities in Chinese and English. The decoding temperature is set to 0.2, and the top-$p$ parameter is set to 0.8.
@@ IV-D Experiment Results
As shown in Table III, on the LibriSpeech test-clean subset, GPT-4o achieved the lowest WER of 3.19%, corresponding to an average reduction of 0.21 percentage points (5.72% relative reduction) and a maximum reduction of 0.34 percentage points (9%). On the test-other subset, a WER of 8.32% was achieved, with a maximum relative reduction of 1.09 percentage points (11.4%).
@@ IV-E Result Analysis
To analyze noun recall within substitution errors, noun filtering was performed using the method proposed in [che-etal-2021-n] for Chinese (AISHELL-1/2) and [bird2009natural] for English (LibriSpeech). The analysis reveals an improvement in noun recall following correction.
@@ V-A Framework Analysis
To evaluate the effectiveness of each framework component, a component-wise ablation study was conducted on AISHELL-1 using DeepSeek-V2. Results are summarized in Table V. Consistent with prior findings [minExploringIntegrationLarge2024], using LLMs with simple prompting for error correction often leads to increased error rates, as reflected by the baseline results.
@@
By integrating iterative reasoning into the framework, the CER was reduced to 4.89%. Finally, implementing answer verification further reduced the CER to 4.69%, effectively mitigating hallucination-induced errors in LLM corrections.
@@ V-C Token Consumption Analysis
To reduce input token consumption during experiments, multiple sentences were grouped together for inference. Empirical observations suggest that this strategy only slightly degrades performance. The specific token consumption is summarized in Table V.